OpenStudy (anonymous):

Two studies were completed in California. One study in northern California involved 1,000 patients; 74% of them experienced flulike symptoms during the month of December. The other study, in southern California, involved 500 patients; 34% of them experienced flulike symptoms during the same month. Which study has the smallest margin of error for a 98% confidence interval?

OpenStudy (anonymous):

The northern California study with a margin of error of 3.2%. The southern California study with a margin of error of 3.2%. The northern California study with a margin of error of 4.9%. The southern California study with a margin of error of 4.9%.

OpenStudy (kropot72):

The margin of error is found from \[\large 2.33\sqrt{\frac{p(1-p)}{n}}\] where p is the sample proportion expressed as a decimal. Therefore the margin of error for the northern California study is \[\large 2.33\sqrt{\frac{0.74\times0.26}{1000}}=you\ can\ calculate.\] Now you need to use the same procedure to find the margin of error for the southern California study.

OpenStudy (anonymous):

so whats the answer?

OpenStudy (anonymous):

@kropot72

OpenStudy (anonymous):

because both A and D are the answer but only one can be chosen!

OpenStudy (anonymous):

ok, A is the correct choice!

OpenStudy (anonymous):

for future users!